### v2

# EasyGen: Easing Multimodal Generation with a Bidirectional Conditional Diffusion Model and LLMs

Abstract
We present EasyGen, an efficient model designed to enhance multimodal understanding and generation by harnessing the capabilities of diffusion models and large language models (LLMs). Unlike existing multimodal models that predominately depend on encoders like CLIP or ImageBind and need ample amounts of training data to bridge modalities, EasyGen leverages BiDiffuser, a bidirectional conditional diffusion model, to foster more efficient modality interactions. EasyGen achieves text generation by training a projection layer linking BiDiffuser and an LLM, and facilities image generation by training an adapter to align the LLM’s text space with the BiDiffuser’s image space. Comprehensive quantitative and qualitative experiments show that EasyGen excels in data-efficient training, high-quality image generation, and extendibility, effectively addressing the challenges in multimodal generation. The source code is available at https://github.com/zxy556677/EasyGen.
EasyGen: Easing Multimodal Generation with a Bidirectional Conditional Diffusion Model and LLMs
Xiangyu Zhao, Bo Liu [^1] Equal contribution., Qijiong Liu [^1] Equal contribution., Guangyuan Shi [^1] Equal contribution., Xiao-Ming Wu Department of Computing, The Hong Kong Polytechnic University {xiang-yu.zhao, bokelvin.liu, jyonn.liu, guang-yuan.shi}@connect.polyu.hk, xiao-ming.wu@polyu.edu.hk
1 Introduction
In recent years, remarkable progress has been made in the field of artificial intelligence generated content (AIGC), notably in technologies like large language models (LLMs) (Chiang et al., 2023; Touvron et al., 2023; Brown et al., 2020; Chowdhery et al., 2022; Zeng et al., 2022) for text generation and diffusion models Rombach et al. (2022); Nichol et al. (2022); Saharia et al. (2022) for visual generation. These breakthroughs have paved the way for the development of multimodal large language models (MLLMs), sparking a recent trend of incorporating extra visual modules into LLMs. Collaborative models, such as Visual ChatGPT (Wu et al., 2023a) and MM-REACT (Yang et al., 2023), strategically use externally pre-trained tools to translate visual information into text descriptions and feed the data into LLMs. However, they are exclusively dependent on pre-trained tools for inference. Contrarily, end-to-end trained models including the BLIP series (Li et al., 2023b), LLaVA series (Liu et al., 2023b, a), MiniGPT-4 (Zhu et al., 2023), and mPLUG-Owl (Ye et al., 2023) focus on mapping image information to the text space of LLMs, enabling LLMs to comprehend visual inputs.
Existing end-to-end models are also not without limitations. First, most of these multimodal models rely on either CLIP (Radford et al., 2021) or ImageBind (Girdhar et al., 2023) as their image encoder. While these encoders excel in learning unified representations that encompass both text and images, they face challenges when it comes to transforming between different modalities. This predicament makes current vision-language models relying heavily on sizable data sets to align CLIP/Bind-encoded images with the language model, due to the disparity between different modalities.
Moreover, the majority of previous multimodal models have concentrated on comprehending multimodal content and lacked the capability to generate multimodal responses, such as content beyond text. Several concurrent works, such as Emu (Sun et al., 2023) and NExT-GPT (Wu et al., 2023), have utilized diffusion models for multimodal generation. Typically, these methods involve training a projection layer to align the output embedding space of the LLM with the input embedding space of the diffusion model (encoded by CLIP’s text encoder) using an MSE loss. However, this approach may lead to the underutilization of the semantic understanding and reasoning capabilities of the LLM, and may introduce information loss in the alignment process, ultimately leading to lower image generation quality compared to the original diffusion model, as elaborated in Sec. 5.6 and Tab. 6.
Figure 1: Our model EasyGen can understand multimodal inputs and generate multimodal responses, as illustrated by model-generated speech bubbles in grey color, which include both text and images.
In this work, we propose EasyGen, an end-to-end model that facilitates multimodal generation with a single bidirectional conditional diffusion model and LLMs, as illustrated in Figure 2. The diffusion model, called BiDiffuser, is obtained by fine-tuning the UniDiffuser (Bao et al., 2023b), with a specific focus on targeted image-to-text and text-to-image tasks. This fine-tuning addresses UniDiffuser’s limitation of attempting to fit all conditional distributions, including those based on noisy inputs, into a single model, which reduces its effectiveness on specific tasks like conditional generation from noise-free inputs. BiDiffuser plays a pivotal role for both text and image generation. In EasyGen, text generation is achieved by training a projection layer that connects BiDiffuser and an LLM, while image generation is facilitated by training an adapter that infuses the text representation of the LLM into BiDiffuser. Figure 1 showcases EasyGen’s ability to handle multimodal inputs and generate appropriate multimodal responses.
EasyGen holds three significant advantages that address the challenges in multimodal generation:
First, EasyGen offers competitive performance in a data-efficient way compared to cutting-edge models, as shown in Tab. 3 (Sec. 5.5). This is due to BiDiffuser’s ability to simplify the alignment of its embedding space with an LLM, which allows for efficient training with less data for image-to-text tasks such as image captioning and VQA.
Second, EasyGen exhibits superior image generation quality, surpassing other end-to-end MLLMs, as shown in Tab. 6 (Sec. 5.6). This is attributed to the adapter’s design (Sec. 4.2), which aligns the LLM’s text space with the diffusion model’s image space, thereby utilizing the LLM’s semantic understanding and reasoning capabilities. In contrast, the projection layers in other MLLMs like NExT-GPT only align the LLM’s text space with the diffusion model’s text space and are not trained by the image denoising objective.
Third, EasyGen can be readily adapted to manage complex vision-language tasks by incorporating more advanced visual encoders or by integrating BiDiffuser into contemporary sophisticated multimodal LLMs like LLaVA to enhance performance, as shown in Tab. 5 (Sec. 5.7).
Figure 2: Overview of EasyGen.
2 Related Work
Multimodal Language Models. Recent research has witnessed a surge of interest in multimodal LLMs, including collaborative models (Wu et al., 2023a; Yang et al., 2023; Shen et al., 2023) and end-to-end methods (Alayrac et al., 2022; Guo et al., 2022; Li et al., 2022; Bao et al., 2021; Wang et al., 2022b, a, a). More recently, some works also explore training LLMs with parameter-efficient tuning (Li et al., 2023b; Zhang et al., 2023a) and instruction tuning (Dai et al., 2023; Liu et al., 2023b; Ye et al., 2023; Zhu et al., 2023; Li et al., 2023a). Different from them, EasyGen is built upon BiDiffuser, which promotes more efficient interactions between modalities.
Multimodal Diffusion Models. Diffusion generative models (Rombach et al., 2022; Ramesh et al., 2021; Nichol et al., 2022; Ruiz et al., 2023) have achieved strong results in text conditioned image generation works. Specifically, Versatile Diffusion (Xu et al., 2023) employs the U-Net (Ronneberger et al., 2015) architecture with a multi-flow design to tackle multiple modalities and tasks, while UniDiffuser (Bao et al., 2023b) adopts the U-ViT (Bao et al., 2023a) framework to treat both image and text as sequential token streams for diffusion calculations. However, these models are unable to complete complex language tasks. EasyGen combines the advantages of diffusion models and LLMs and achieves competitive performance in both image-to-text and text-to-image tasks.
Multimodal Response Generation. Recent research has made significant advancements in multimodal response generation (Koh et al., 2023b; Tang et al., 2023; Zhang et al., 2023b; Wu et al., 2023b; Pan et al., 2023; Koh et al., 2023a; Sun et al., 2023; Dong et al., 2023) using text-to-image models such as Stable Diffusion. However, the lack of semantic understanding capability in the CLIP text encoder may result in low-quality generated images. EasyGen addresses this issue by transferring knowledge from LLM to BiDiffuser via an adapter, enabling the creation of high-quality textual semantic representations for text-to-image generation.
3 Basics of Diffusion Models
Unconditional Generation. Given a data sample taken from a real data distribution $\mathbf{x}_{0}\sim q(\mathbf{x}_{0})$, diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) first destruct the data by constructing a Markov forward process and gradually injecting noise to the data:
$q(\mathbf{x}_{1:T}|\mathbf{x}_{0})$ $=\prod_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1}),$ (1)
$\quad q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$ $=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I}),$
where $\beta_{t}\in(0,1)$ is the variance added at diffusion step $t$. Then, they learn to reverse the process:
$p(\mathbf{x}_{0:T})$ $=p(\mathbf{x}_{T})\prod_{t=1}^{T}p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t}),$ (2)
$\quad p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ $=\mathcal{N}(\mathbf{x}_{t-1};\mu_{t}(\mathbf{x}_{t},t),\sigma_{t}^{2}\mathbf{I}),$
where $p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{T};0,\mathbf{I})$ is the standard Gaussian distribution and $\mu_{t}(\cdot)$ is the parameterization of the predicted mean. Diffusion models are trained to maximize the marginal likelihood of the data $\mathbb{E}[\log p_{\theta}(\mathbf{x}_{0})]$, and the canonical objective is the variational lower bound of $\log p_{\theta}(\mathbf{x}_{0})$. Denoising diffusion probabilistic models (Ho et al., 2020) generate samples $\mathbf{x}_{t}\sim q(\mathbf{x}_{t}|\mathbf{x}_{0})$ by injecting noise $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$ to the data $\mathbf{x}_{0}$, and train a network $\bm{\epsilon}_{\theta}(\cdot)$ to predict the added noise $\bm{\epsilon}$ using a standard mean squared error loss:
$\mathcal{L}:=\mathbb{E}_{\mathbf{x}_{0},{\bm{\epsilon}},t}[\|{\bm{\epsilon}}-\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)\|^{2}].$ (3)
Conditional Generation.
For conditional generation, a paired data $(\mathbf{x}_{0},\mathbf{y}_{0})$ is given, and the aim is to model the conditional data distribution $q(\mathbf{x}_{0}|\mathbf{y}_{0})$, where $\mathbf{y}_{0}$ can be image class or text prompt. Conditional generation includes classifier guidance (Dhariwal and Nichol, 2021) and classifier-free guidance (Ho and Salimans, 2021). Classifier guidance requires training an extra classifier on noisy data at inference time to improve sample quality. For classifier-free guidance, no classifier needs to be trained. The denosing network $\bm{\epsilon}_{\theta}(\mathbf{x}_{t}|\mathbf{y}_{0})$ simply conditions on the information encoded in $\mathbf{y}_{0}$. At inference time, with a guidance scale $s$, the modified score estimate is further in the direction of $\bm{\epsilon}_{\theta}(\mathbf{x}_{t}|\mathbf{y}_{0})$ and away from the unconditional model $\bm{\epsilon}_{\theta}(\mathbf{x}_{t}|\emptyset)$ ($\emptyset$ is a null token):
$\hat{\bm{\epsilon}}_{\theta}(\mathbf{x}_{t}|\mathbf{y}_{0})=\bm{\epsilon}_{\theta}(\mathbf{x}_{t}|\emptyset)+s\cdot(\bm{\epsilon}_{\theta}(\mathbf{x}_{t}|\mathbf{y}_{0})-\bm{\epsilon}_{\theta}(\mathbf{x}_{t}|\emptyset)).$
4 Proposed Model: EasyGen
We propose EasyGen, a model capable of processing multimodal inputs and generating multimodal outputs. It achieves easy multimodal generation by leveraging a bidirectional conditional diffusion model to effectively bridge the gap between different modalities and an LLM to comprehend multimodal tasks and produce textual responses containing cues for multimodal message creation. In the subsequent section, we outline the multimodal generation process of EasyGen.
4.1 Pre-training BiDiffuser: A Bidirectional Conditional Diffusion Model
Since the text space of LLMs is discrete, to minimize the disparity between the output of a diffusion model and the input of LLMs, we leverage Unidiffuser, a unified diffusion model capable of transforming images into the discrete text space. During the training process, UniDiffuser injects noise $\bm{\epsilon}^{x}$ and $\bm{\epsilon}^{y}$ to a set of paired image-text data $(\mathbf{x}_{0},\mathbf{y}_{0})$ and generates noisy data $\mathbf{x}_{t^{x}}$ and $\mathbf{y}_{t^{y}}$, where $0\leqslant t^{x},t^{y}\leqslant T$ represent two individual timesteps (perturbation levels). It then trains a joint denoising transformer U-ViT (Bao et al., 2023a) $\bm{\epsilon}_{\theta}(\mathbf{x}_{t^{x}},\mathbf{y}_{t^{y}},t^{x},t^{y})$ to predict the noise $\bm{\epsilon}^{x}$ and $\bm{\epsilon}^{y}$ by minimizing the mean squared error loss:
$\mathbb{E}_{\bm{\epsilon}^{x},\bm{\epsilon}^{y},\mathbf{x}_{0},\mathbf{y}_{0}}[\|[\bm{\epsilon}^{x},\bm{\epsilon}^{y}]-\bm{\epsilon}_{\theta}(\mathbf{x}_{t^{x}},\mathbf{y}_{t^{y}},t^{x},t^{y})\|^{2}],$
where the output of $\bm{\epsilon}_{\theta}$ is the concatenation of the estimated noise $\bm{\epsilon}_{\theta}^{x}$ and $\bm{\epsilon}_{\theta}^{y}$, i.e., $\bm{\epsilon}_{\theta}=[\bm{\epsilon}_{\theta}^{x},\bm{\epsilon}_{\theta}^{y}]$.
By predicting $\bm{\epsilon}_{\theta}(\mathbf{x}_{t^{x}},\mathbf{y}_{t^{y}},t^{x},t^{y})$ for any $t^{x}$ and $t^{y}$, UniDiffuser learns all distributions related to $(\mathbf{x}_{0},\mathbf{y}_{0})$ simultaneously. This includes all conditional distributions: $q(\mathbf{x}_{0}|\mathbf{y}_{0})$ for text-to-image generation, $q(\mathbf{y}_{0}|\mathbf{x}_{0})$ for image-to-text generation, and those conditioned on noisy input, i.e., $q(\mathbf{x}_{0}|\mathbf{y}_{t^{y}})$ and $q(\mathbf{y}_{0}|\mathbf{x}_{t^{x}})$, for $0<t^{x},t^{y}\leq T$. Learning a conditional distribution $q(\mathbf{x}_{0}|\mathbf{y}_{t^{y}})$ or $q(\mathbf{y}_{0}|\mathbf{x}_{t^{x}})$ can be seen as learning a distinct task. From a multitask learning perspective, due to limited network capacity, learning many tasks simultaneously (i.e., fitting all distributions to a single network) may result in task competition or task conflict, ultimately leading to suboptimal performance in particular tasks such as $q(\mathbf{x}_{0}|\mathbf{y}_{0})$ and $q(\mathbf{y}_{0}|\mathbf{x}_{0})$.
Figure 3: The training of BiDiffuser involves finetuning the denoising transformer U-ViT in UniDiffuser with a joint objective of image-to-text and text-to-image tasks.
To resolve this issue and enhance the performance of both image-to-text and text-to-image generation tasks, we finetune UniDiffuser with exclusive emphasis on the two tasks:
$\mathcal{L}_{d}=\mathbb{E}_{\bm{\epsilon}^{x},\bm{\epsilon}^{y},\mathbf{x}_{0},\mathbf{y}_{0}}[\|\bm{\epsilon}^{x}-\bm{\epsilon}_{\theta}^{x}(\mathbf{x}_{t^{x}},\mathbf{y}_{0},t^{x},0)\|^{2}+$
$\alpha\|\bm{\epsilon}^{y}-\bm{\epsilon}_{\theta}^{y}(\mathbf{x}_{0},\mathbf{y}_{t^{y}},0,t^{y})\|^{2}].$
where $\alpha$ is a hyperparameter to balance the learning paces of the two tasks. As depicted in Figure 3, our training objective entails predicting the text $\mathbf{y}_{0}$ based on the input image $\mathbf{x}_{0}$ and vice versa, where the input conditions for the model are noise-free. We name the finetuned model “BiDiffuser”, signifying its specialized ability in bidirectional conditional generation.
4.2 Pre-training an Adapter to Enhance BiDiffuser’s SUR Capability
BiDiffuser uses the text encoder of CLIP, which is trained with image-text contrastive learning, limiting its semantic understanding and reasoning (SUR) ability for image generation. Drawing inspiration from Zhong et al. (2023), we utilize LLMs to enhance the SUR capability of LLMs. Specifically, we design an adapter that employs the attention mechanism to integrate the semantic information from LLM’s last hidden state $f_{\text{LLM}}(\cdot)$ into the CLIP text encoder $f_{\text{CLIP}}(\cdot)$. The adapter consists of a projection layer $\text{MLP}(\cdot)$ and a cross-attention layer $\text{Att}(\cdot)$. Given a paired image-text data $(\mathbf{x}_{0},\mathbf{y}_{0})$, we can get $y_{\text{sur}}$ with enhanced SUR via the adapter:
$y_{\text{sur}}=\text{Att}(f_{\text{CLIP}}(\mathbf{y}_{0})W^{Q},\text{MLP}(f_{\text{LLM}}(\mathbf{y}_{0}))W^{K},$
$\text{MLP}(f_{\text{LLM}}(\mathbf{y}_{0}))W^{V}).$
Then, the semantic input to BiDiffuser is the combination of $y_{\text{sur}}$ and the CLIP text encoding of $\mathbf{y}_{0}$:
$y_{0}=\lambda\cdot y_{\text{sur}}+(1-\lambda)\cdot f_{\text{CLIP}}(\mathbf{y}_{0}),$
where $\lambda$ is a balancing parameter. We train the adapter by freezing BiDiffuser and minimizing
$\mathcal{L}_{\text{ada}}=\mathbb{E}_{\bm{\epsilon}^{y},\mathbf{x}_{0}}[\|\bm{\epsilon}^{x}-\bm{\epsilon}_{\theta}^{x}(\mathbf{x}_{t^{x}},y_{0},t^{x})\|^{2}],$ (4)
where ${\epsilon}_{\theta}^{x}$ is not updated as BiDiffuser is frozen.
4.3 Image-to-Text Generation
BiDiffuser can convert images into vectors in the text space, facilitating alignment with the vector space of LLMs. In the following, we show how BiDiffuser can be integrated with LLMs to perform image-to-text generation tasks such as image captioning and visual question answering (VQA).
4.3.1 Aligning BiDiffuser with LLMs
We connect BiDiffuser and LLMs via a simple projection layer, which maps text embeddings obtained from the output of the diffusion model to the embedding space of LLMs. As shown in Figure 4, the alignment can take place either prior to the LLM (Pre-Align manner) or between its encoder and decoder components (Mid-Align manner).
(a) Pre-Align manner.
(b) Mid-Align manner.
Figure 4: Two different ways of aligning BiDiffuser with LLMs.
Pre-Align Manner. As shown in Figure 3(a), the projection layer is placed before the LLM to map the output of BiDiffuser (image representations) to the text embedding space of the LLM. The text embedding of the input image is then concatenated with the embeddings of the textual instructions and fed to the LLM for decoding. To synchronize the text space of BiDiffuser with that of the LLM, we propose to use the image-grounded text generation (ITG) objective to drive the model to generate texts based on the input image by computing the auto-regressive loss:
$\mathcal{L}_{\text{ITG}}=-\frac{1}{L}\sum_{l=1}^{L}\log p_{\phi}(w_{l}^{g}|w_{<l}^{g},I,T_{I}),$ (5)
where $w^{g}=(w_{1}^{g},...,w_{L}^{g})$ represents the ground-truth caption of image $I$ with length $L$, $T_{I}$ is the text instruction, and $\phi$ denotes the model parameters, which include the parameters of the projection layer and the LLM.
Mid-Align Manner. As shown in Figure 3(b), the projection layer is placed between the LLM’s encoder and decoder, aiming to map the output of BiDiffuser to the embedding space of the text that is encoded by the LLM’s encoder. Particularly, we argue that the output of BiDiffuser, once mapped by the projection layer and denoted as $\mathbf{d}_{\mathrm{diff}}$, should align with the image caption that is encoded by the LLM’s encoder, denoted as $\mathbf{d}_{\mathrm{llm}}$. Therefore, to accurately learn the alignment between the image and text representations, in addition to the ITG loss in Eq. 5, we also employ an image-text distance minimization (ITDM) loss:
$\mathcal{L}_{\text{ITDM}}$ $=\frac{1}{N}\sum_{i=1}^{N}\|\mathbf{d}_{\mathrm{diff}}-\mathbf{d}_{\mathrm{llm}}\|_{2}^{2},$ (6)
$\mathcal{L}_{\text{mid}}$ $=\mathcal{L}_{\text{ITG}}+\mathcal{L}_{\text{ITM}}.$
where $N$ is the batch size, and $\mathcal{L}_{\text{mid}}$ is the overall loss. In this manner, the model parameters $\theta$ only include the parameters of the projection layer.
After the alignment, EasyGen gains the capability of zero-shot image-to-text generation, including tasks such as image captioning and VQA.
4.3.2 Instruction-Tuning LLMs
When aligning BiDiffuser with an LLM, we perform instruction-tuning on the LLM to equip it with the capability of understanding multimodal tasks. We designed different instructions for different LLMs, as shown in Table 11. General instruction template is denoted as follows:
USER: <Img><image></Img> + Instruction. Assistant: <answer>.
For the <image> placeholder, we substitute it with the output of BiDiffuser. To avoid over fitting to the specific task and counter the model’s inclination to generate excessively short outputs, we have devised specific instructions (see Table 10), which enable the LLM to produce concise responses when necessary. For different tasks, the distinct instruction templates are as outlined in Appendix F.
4.4 Text-to-Image Response Generation
Most of existing multimodal models, including the BLIP series and LLaVA series are unable to provide a multimodal response as they are primarily designed to generate only textual outputs. On the other hand, Emu (Sun et al., 2023) takes a unified approach to predict the subsequent visual or textual token in an auto-regressive manner, but it is heavily reliant on vast quantities of training data. Contrary to the limitations of these existing models, EasyGen, by leveraging the bidirectional generation capability of BiDiffuser and the inference capability of LLMs, can produce accurate and high-quality visual response with ease.
To tackle multimodal response generation tasks such as PhotoChat (Zang et al., 2021), we first finetune the LLM to generate detailed image captions based on dialogue context. Then, we employ BiDiffuser to create the corresponding images with the produced captions. Specifically, we replace the image featured in the dialogue with its corresponding descriptive caption, encapsulating it with task-specific tokens <Img>,</Img> and constructing the following instruction templates:
USER: Dialog history. Assistant: <response> + <Img><caption></Img>.
Figure 5: Text-to-image generation by EasyGen. LLM generates the response and description of the image. BiDiffuser generates images based on the description.
When <caption> appears in response, it represents the generated description of the image. So we can use LLM’s original auto-regressive training objective. Specifically, we compute the probability of the target caption by:
$\mathcal{L}_{t2t}=-\frac{1}{L}\sum_{l=1}^{L}\log p_{\vartheta}(w_{l}^{c}|w_{<l}^{c},H),$ (7)
where $w^{c}=(w_{1}^{c},...,w_{L}^{c})$ represents the caption of image $x_{0}$ with length $L$, $H$ is the dialogue history, and $\vartheta$ denotes the LLM’s parameters. Considering the potential for alignment deviation in text alone, we utilize the output of the LLM as the conditional component of the diffusion model. This is achieved through the adapter, which directly contributes to the denoising process. The loss function for the denoising process of a noisy image $x_{t^{x}}$ is formulated in a way that is similar to Eq. 4:
$\mathcal{L}_{t2i}$ $=\mathbb{E}_{\bm{\epsilon}^{y},\mathbf{x}_{0}}[\|\bm{\epsilon}^{x}-\bm{\epsilon}_{\theta}^{x}(\mathbf{x}_{t^{x}},y_{0},t^{x})\|^{2}],$ (8)
where ${\epsilon}_{\theta}^{x}$ is not updated and we only train the parameters of LLM and adapter. The overall loss for text-to-image task is:
$\mathcal{L}_{all}$ $=\mathcal{L}_{t2i}+\mathcal{L}_{t2t}.$ (9)
Training with the instruction data enables our model to not only produce text responses but also perform image intent classification and generate image captions that BiDiffuser can interpret.
5 Experiments
5.1 Experimental Setup
We initialize encoder-decoder LLM from FlanT5-XL or decoder-only LLM from Vicuna-7B, along with the utilization of the diffusion module from BiDiffuser. During the alignment process, we maintain the frozen state of the BiDiffuser. The statistics of the datasets for pre-training, alignment and instruction-tuning can be found in Appendix B. For the image captioning task, EasyGen is evaluated on both the MS-COCO (Lin et al., 2014) Karpathy test set and the NoCaps (Agrawal et al., 2019) validation set. For the VQA task, we evaluated on OK-VQA (Marino et al., 2019) validation set and GQA (Hudson and Manning, 2019) test-dev set.
To adapt the model for multimodal dialogue generation, we fine-tune the LLM and projection layers on the PhotoChat dataset. We incorporate photo-sharing activities into the dialogue context by generating <Img><caption></Img>, and utilize cross-entropy loss exclusively for fine-tuning the multimodal generation task. Given the limited expressiveness of image descriptions in the PhotoChat dataset, as evidenced by Figure 7 in Appendix I, we regenerate image annotations in a text format similar to that used in MS-COCO.
Model Dataset Size NoCaps (val) COCO (Karpathy) OK-VQA GQA
PT IT CIDEr SPICE BLEU@4 CIDEr Accuracy Accuracy
BLIP (Li et al., 2022) 129M - 113.2 14.8 40.4 136.7 - -
Flamingo (Alayrac et al., 2022) 1.8B - - - - 138.1 50.6 -
BLIP-2 OPT-6.7B (Li et al., 2023b) 129M - 121.0 15.3 43.5 145.2 36.4 36.4
BLIP-2 FlanT5XL (Li et al., 2023b) 129M - 121.6 15.8 42.4 144.5 39.4 44.4
InstructBlip 7B (Dai et al., 2023) 129M 1.2M 123.1 - 40.8 140.7 $61.0^{\star}$ $49.2^{\star}$
MiniGPT-4 (Zhu et al., 2023) - 5M 42.4 - - - 37.5 30.8
LLaVA (Liu et al., 2023b) 558K 158K 33.1 - 7.9 30.0 54.4 41.3
EasyGen FlanT5XL 169K 90K 121.2 15.5 43.5 145.7 41.1 37.2
EasyGen Vicuna-7B 169K 90K 121.8 15.8 42.4 144.6 45.2 44.6
Table 1: Evaluations of EasyGen and baselines on various image understanding tasks. PT, IT indicate sample sizes in the pretraining and instruction tuning stages respectively. EasyGen’s results on NoCaps, OK-VQA and GQA were obtained in a zero-shot setting. $\star$ denotes that the model was trained on other VQA datasets.
Model Response Generation Image
[…]
Divter Sun et al. 6.5/1.7 59.6 5.69 29.16
Maria Liang et al. 13.8/9.2 48.7 15.17 -
EasyGen FlanT5 22.3/18.7 13.3 17.24 10.30
EasyGen Vicuan 23.6/19.9 11.3 18.85 9.72
+ w/o adapter - - - 10.16
Table 2: Evaluation on the PhotoChat dataset.
MLLM Sample Size Cosine Similarity $\uparrow$ MSE $\downarrow$ MiniGPT-4 5M 0.0016 6.2031 LLaVA v1.5 558K -0.0026 0.8433 Emu 2B 0.0054 0.4062 EasyGen 169K 0.0128 0.0338 Table 3: Data efficiency. Avg. Cosine similarity and mean square error between the projected representations and their respective captions embedded by LLM.
LLM Diffusion Model Alignment NoCaps COCO(Karpathy) OK-VQA
CIDEr SPICE BLEU@4 CIDEr Accuracy
[Uncaptioned image] T5 UniDiffuser Pre-Align 62.4 18.0 26.8 90.7 33.0
[Uncaptioned image] T5 BiDiffuser Pre-Align 119.1 25.5 42.6 145.1 41.1
[Uncaptioned image] T5 BiDiffuser Mid-Align 121.2 25.1 43.5 145.7 31.5
[Uncaptioned image] T5 BiDiffuser Mid-Align 121.5 25.3 43.6 145.7 36.4
[Uncaptioned image] Vicuna-7B BiDiffuser Pre-Align 121.8 24.9 42.4 144.6 45.2
[Uncaptioned image] Vicuna-7B BiDiffuser Pre-Align 119.0 24.6 40.3 140.3 42.7
Table 4: Ablation study on image captioning and VQA tasks. [Uncaptioned image] / [Uncaptioned image] denotes tuning/freezing the LLM.
5.2 Evaluation
We evaluate EasyGen on various vision-language tasks including image captioning (MS-COCO (Lin et al., 2014), NoCaps (Agrawal et al., 2019)), visual question answering (OK-VQA (Marino et al., 2019), GQA (Hudson and Manning, 2019)), and multimodal dialog generation (PhotoChat (Zang et al., 2021)). We use BLIP (Li et al., 2022), Flamingo (Alayrac et al., 2022), BLIP-2 (Li et al., 2023b), InstructBlip (Dai et al., 2023), MiniGPT-4 (Zhu et al., 2023), and LLaVA (Liu et al., 2023b) as baselines for image-to-text tasks, and Maria (Liang et al., 2021) and Divter (Sun et al., 2021) as baselines for the multimodal response generation task. See details in Appendix C and  E.
5.3 Overall Results
Tab. 1 presents the evaluation results for each baseline and our models on MS-COCO and VQA (zero-shot) datasets. EasyGen outperforms most of the baseline models on both the COCO test set and NoCaps validation set (zero-shot transfer). Despite being pre-trained on a small dataset (MS-COCO), EasyGen’s performance on the image captioning generation task is comparable to models trained on larger datasets. Additionally, on the OK-VQA and GQA datasets, EasyGen demonstrates improved performance compared to other models of a similar scale, achieving higher accuracy even with a simple greedy search decoding method.
In Tab. 2, the evaluation results on the PhotoChat dataset are presented. Our method exhibits clear advantages in terms of PPL, indicating strong performance on response generation task. Because of the image descriptions in the PhotoChat dataset are overly concise, we utilized EasyGen to regenerate the image descriptions, which improved the performance of our model on image generation compared to other models. Additionally, with the adapter, EasyGen is capable of generating images of superior quality.
5.4 Ablation Study
In Tab. 4, we examine the impact of freezing/tuning BiDiffuser and the LLM. It can be observed that frozen Mid-Align method outperforms Pre-Align method in image captioning, which shows ITDM loss function is effective. However, the frozen Mid-Align method exhibits inferior performance in the VQA task. We hypothesize that this is due to the integration of mid-aligned target image features with query information, and the projection layer is insensitive to instruction information. We conduct instruction-tuning on Pre-Align T5 and Vicuna. Compared to models at the same scale, these instruction-tuned models achieve superior results.
Model IT VQAv2 (test-dev) TextVQA MMBench (test)
MiniGPT-4 (Zhu et al., 2023) 5M - 19.4 23.0
InstructBLIP Vicuna-7B (Dai et al., 2023) 1.2M - 50.1 33.9
LLaVA-1.5 Vicuna-7B (Liu et al., 2023a) 665K 78.5 58.2 65.2
LLaVA-1.5 Vicuna-13B (Liu et al., 2023a) 665K 80.0 61.3 67.8
EasyGen Vicuna-7B w/ ViT-L 251K 79.4 57.9 63.9
LLaVA-1.5 Vicuna-7B w/ EasyGen 665K 80.2 58.8 66.1
LLaVA-1.5 Vicuna-13B w/ EasyGen 665K 80.5 61.5 69.2
Table 5: Evaluation of EasyGen variants and baselines on more complex VQA tasks and the latest MMBench benchmark. “w/ EasyGen” means incorporating the core components of our model into existing models as depicted in Figure 6 in Appendix F. EasyGen variants rank among the top models on the leaderboard of MMBench.
5.5 Data Efficiency in Training
In Tab. 3, we examine the data efficiency of different image encoders for alignment with LLMs. EasyGen uses BiDiffuser, which maps images to the text space, simplifying alignment with LLMs. To assess the quality of visual representations, we measured the distance between the projected representations and their respective captions embedded by an LLM. We randomly selected 1,000 images with their corresponding captions from the MSCOCO dataset. The results show that our model, EasyGen, aligns significantly better with the LLM compared to other CLIP-based MLLMs, despite using less data for alignment. This indicates the effectiveness of our approach in achieving strong alignment with LLMs.
MLLM FID $\downarrow$ Diffusion Model FID $\downarrow$ NExT-GPT 11.28 (+0.07) SD 11.21 Emu 11.66 (+1.73) SD v1.5 9.93 Fine-tuned on MS-COCO EasyGen 7.68 (-0.44) UniDiffuser 8.12 + w/o adapter 7.89 (-0.23) UniDiffuser 8.12 Fine-tuned on original dataset (zero-shot) EasyGen 9.16 (-0.55) UniDiffuser 9.71 + w/o adapter 9.52 (-0.19) UniDiffuser 9.71 Table 6: Comparing the image generation quality of end-to-end MLLMs and their corresponding diffusion models on the MS-COCO validation set (256 × 256). Our EasyGen surpasses the original diffusion model, while other MLLMs fall short in comparison.
5.6 Image Generation Quality
Tab. 6 evaluates the generated image’s quality of MLLMs on MS-COCO validation set, using 30K randomly selected prompts to compute the FID score on generated images. To confirm the efficacy of our approach, we fine-tuned our method on a portion of the original data (LIAON-COCO) and the MS-COCO train set, respectively. While other models resulted in a decrease in image generation performance compared to the corresponding diffusion model, EasyGen outperformed the original UniDiffuser due to the fine-tuned BiDiffuser and the adapter module. This suggests that the alignment between LLM and diffusion model’s text space used in other MLLMs may be unnecessary. Instead, generating images from texts/captions and aligning them with the diffusion model’s image space appears to be an effective solution.
5.7 Extendability
Tab. 5 explores the extensibility of our method from two aspects. Firstly, we aim to enhance the performance of EasyGen on complex tasks such as VQA and OCR by integrating more powerful visual encoders. Considering the potential information dilution or omission when using BiDiffuser to convert images to text space, we choose to integrate CLIP ViT-L/14 as the image encoder (as depicted in Figure 6 in the Appendix). During this process, we freeze CLIP and BiDiffuser while fine-tuning the parameters of the LLM and projection layers. The results presented in Tab. 5 include performance on traditional short QA and the modern benchmark MMBench Liu et al. (2023c). With CLIP ViT-L, EasyGen’s performance is better than LLaVA on the VQAv2 dataset, demonstrating that BiDiffuser can effectively assist LLM in understanding images. Secondly, we investigate the plug-and-play capability of BiDiffuser, as it can also be integrated into other MLLMs (with the same LLMs) to improve their performance. The results demonstrate that with BiDiffuser, LLaVA-1.5 could achieve better performance. We speculate that BiDiffuser provides guidance information to MLLMs, enabling them to better understand the details of CLIP encoded images.
6 Conclusion
We have introduced EasyGen, a model that facilitates multimodal understanding and generation. Compared to existing models, EasyGen offers a more efficient solution by employing BiDiffuser, a bidirectional diffusion model. This allows for more effective modal interactions, handling both image-to-text and text-to-image generations by the fusion of BiDiffuser and LLMs. Additionally, EasyGen can be easily integrated into existing advanced multimodal LLMs to further boost their performance.
7 Limitations
This section aims to highlight the limitations of our work and provide further insights into the research in this area. Our model relies on diffusion for multi-modal interaction, which means that the text-to-image and image-to-text processes may take longer. In our experiments, we tested the performance of our model on one A100 (80G) GPU. During inference, using 1000 image-caption pairs, EasyGen took approximately 2.95 seconds for the caption generation task (with the diffusion module taking about 2.41 seconds) and around 4.96 seconds to generate an image. We believe it would be beneficial to explore more efficient sampling methods, such as DPM-Solver++ Lu et al. (2022), to improve the overall efficiency of EasyGen.
Appendix A Ethics Statement
We adhere to the ACL Ethics Policy and have conducted our research using publicly available repositories and datasets. Our primary focus is on investigating the integration of diffusion models and LLMs for multimodal generation. Therefore, the results should be seen as AI-generated content. While we have not observed deliberate harmful content, the model has the potential to generate such content if triggered. We have taken steps to minimize this risk through fine-tuning on public datasets, but caution is still exercised. In future, we will prioritize improving downstream performance and exploring methods to enhance control over the generation process. To ensure reproducibility and support future research, we have made all resources publicly available and provided proper citations to previous research within the code.
Appendix B Datasets
We test the effectiveness of EasyGen by experimenting on different tasks including image captioning, visual question answering (VQA), and multimodal dialogue tasks. Table 7 shows the statistics of the pre-training datasets for BiDiffuser, alignment and VQA tasks.
We use the MS-COCO (Lin et al., 2014) dataset for image captioning. Following BLIP-2 (Li et al., 2023b), we fine-tune EasyGen on MS-COCO and evaluate its performance on the Karpathy test set and the NoCaps (Agrawal et al., 2019) validation set. In MS-COCO, each image typically has five captions that convey similar meanings. The training set consists of 82,783 images with 414,113 captions, while the COCO Karpathy test set has 5,000 images and the NoCaps validation set has 4,500 images.
For multimodal dialogue, we utilize the PhotoChat (Zang et al., 2021) dataset, which is a high-quality dataset consisting of 10,917 images and 12,286 dialogues. Each dialogue is associated with a user image and its corresponding text description. The dataset is divided into 10,286 training instances, 1,000 development instances, and 1,000 testing instances. Moreover, PhotoChat includes photo-sharing activities, defined as the process of creating <Img><caption></Img> in this study. Each conversation in PhotoChat is broken down and constructed into multiple samples so that each round of responses can be learned. Specifically, we regard the first three turns as the dialog context, and the subsequent turns as the prediction targets. By converting the dialogues of this dataset into the form mentioned in 4.4, we obtained 49,240 train, 4,792 dev, and 4,836 test dialogue pairs.
For the VQA task, we conduct a quantitative evaluation on both the OK-VQA (Marino et al., 2019) validation set (5,046 questions) and the GQA (Hudson and Manning, 2019) test-dev set (12,578 questions). As shown in Table 4, for the frozen LLM, following BLIP-2, we employ the length penalty in beam search to encourage short answer generation. On the contrary, for the tuned LLM, we use the VQA instructions (as shown in Table 9) to do instruction tuning during the alignment process. The data for instruction tuning is constructed by randomly selecting 5K data from VQAv2 (Goyal et al., 2017) and 5K data from Visual Dialog (Murahari et al., 2019) training set.
Data types Dataset Size BiDiffuser Alignment Fine-tuning
Caption MS-COCO caption (Lin et al., 2014) 83K
Visual Genome (Krishna et al., 2017) 86K
Multimodal instruction LLaVA dataset Liu et al. (2023b) 80K
VQA VQAv2 (Goyal et al., 2017) 83K -
AOK-VQA (Schwenk et al., 2022) 66K
OCR-related tasks Text Captions (Sidorov et al., 2020) 22K
TextVQA (Singh et al., 2019)
Table 7: Description of datasets used in our alignment and VQA fine-tuning stages. Noting that in alignment process, we used 5K images from VQAv2 dataset.
Dataset Task Split Metric
Image-to-Text MS-COCO (Lin et al., 2014) Image captioning Test CIDEr, BLEU, SPICE
NoCaps (Agrawal et al., 2019) Image captioning Val CIDEr, SPICE
OK-VQA (Marino et al., 2019) VQA Val Accuracy
GQA (Hudson and Manning, 2019) VQA Test Accuracy
Multimodal Generation PhotoChat Zang et al.,2021 Image dialogue Test PPL, BLEU, ROUGE, FID
Table 8: Summary of the evaluation datasets and metrics.
Task Instruction Template
Image Captioning USER: <image>+random[query] Assistant:
LLaVA 80K USER: Please answer question from this image: <image> Question: <question> Assistant:
USER: Image: <image> Question: <question> Assistant:
USER: Answer question <question> through the image <image> Assistant:
Multimodal Dialogue USER: Dialog history+<photo>+Dialogue history Assistant:
VQA USER: Image: <image> Question: <question> Short answer: Assistant:
USER: Image: <image> Question: <question> Answer the option’s letter. Assistant:
Table 9: Examples of task instruction templates. <image> represents the input image, <question> denotes the question in the VQA and LLaVA 80K dataset, and <photo> is the image description of the input image.
Appendix C Baselines
We compare our proposed model with the following state-of-the-art baselines:
BLIP (Li et al., 2022) is a multimodal mixture of encoder-decoder. It can be used as an image-based text encoder or decoder. We use it to perform caption generation and VQA tasks.
BLIP-2 (Li et al., 2023b) is pre-trained through bootstrapped learning from frozen visual encoder and LLMs using an efficient pre-training strategy.
Flamingo (Alayrac et al., 2022) incorporates new cross-attention layers into Chinchilla language model (Hoffmann et al., 2022) to inject visual features, and pre-trains the new layers on billions of image-text pairs. We use it to perform caption generation and VQA tasks.
InstructBlip (Dai et al., 2023) is a vision-language instruction tuning framework that is trained with BLIP-2 and capable of solving various visual language tasks.
MiniGPT-4 (Zhu et al., 2023) utilizes a single projection layer to align visual information from a pre-trained vision encoder with an LLM. It employed the same visual encoder as used in BLIP-2.
LLaVA (Liu et al., 2023b) employs a solitary projection layer to convert image features extracted from the pre-trained CLIP-ViT-L/14 visual encoder into the language embedding space of Vicuna.
Maria (Liang et al., 2021) is a neural conversation agent which can leverage visual world experiences sourced from a vast image index. It possesses the ability to fetch a relevant image specific to the conversation and extract visual knowledge from it.
Divter (Sun et al., 2021) focuses on exploring multimodal dialogue generative models. Given the dialogue context, this model first generates a text response or image description and then generates an image according to the description.
Figure 6: Model’s architecture for VQA finetuning. The module with blue background is referred to as BiDiffuser, while the rest is the architecture of MLLM using CLIP as the image encoder (such as LLaVA).
Appendix D Evaluation
For evaluating the quality of text generation, we utilize metrics such as BLEU, Rouge-L, Accuracy, and PPL (Perplexity). Additionally, following the approach of Vicuna (Chiang et al., 2023) and LLaVA (Liu et al., 2023b), we employ ChatGPT to assess the generated responses from our model. Specifically, for the image captioning task, we randomly select 30 images from the MS-COCO Karpathy split and then let ChatGPT score the responses generated by EasyGen and the baseline models. ChatGPT evaluates the models’ responses based on relevance, details, and accuracy and assigns an overall score between 1 and 10, with a higher score indicating better performance. To evaluate the quality of image generation, we use the Frechet Inception Distance (FID) score (Heusel et al., 2017), which measures the divergence between two multivariate normal distributions.
Appendix E Implementation Details
LLM
During the alignment process, we utilize the AdamW optimizer with $\beta_{0}$ = 0.9, $\beta_{1}$ = 0.99, and weight decay of 0. The LLMs are trained with a cosine learning rate of 2e-5 and a warmup rate of 0.03. We use a batch size of 96 for the frozen LLMs and 32 for the tuned LLMs. During training, we convert the LLMs (FlanT5XL/Vicuna-7B) to BFloat16/FP16 and BiDiffuser to FP16. During the VQA tuning process, we use CLIP ViT-L/14 336px as additional image encoder. We finetune EasyGen on mixture datasets for 1 epoch with a batch size of 32. We adopt the AdamW optimizer with $\beta$ = (0.9, 0.99) with the learning rate is 2e-5. We use a cosine learning rate decay with a learning rate is 2e-5 and warmup ration is 0.03.
Diffusion Module
We inherit the settings from UniDiffuser and utilize pre-trained weights from its checkpoint for our text-to-image generator. The model is fine-tuned on the MS-COCO and VG dataset, which contains images with a resolution of $512\times 512$, for 10 epochs with a batch size of 312. For all of our sampling processes, we employ DPM-Solver with 50 steps.
1. Describe the image concisely.
2. Provide a brief description of the given image.
3. Can you describe this image briefly?
4. Provide a summary of visual elements depicted in the image.
5. Give me the essential characteristics of the photograph in a
concise manner.
6. Rephrase the image depicted in a concise manner.
7. Describe the objects in this image no in detail.
8. Please introduce the image for me briefly.
9. Give me the image’s short descriptions.
10. Please provide a general depiction of the image presented.
Table 10: For the image captioning task, a query instruction is randomly selected.
Appendix F Instruction Tuning
We list the instructions for different tasks in the main paper in Table 9. Specifically, the queries used to describe image contents are presented in Table 10. Table 9 shows the templates used in Vicuna, if the LLM is FlanT5, kindly use “Human” to substitute “USER” in the instruction templates. Model architecture for VQA finetuning is shown in Figure 6. EasyGen integrates the outputs of BiDiffuser with images encoded by CLIP ViT-L/14. We freeze CLIP and BiDiffuser while only tuning the parameters of the LLM and projection layers.
Different Instruction Templates: [Uncaptioned image] Caption Generation Provide a brief description of the given image. Assistant: Response Buses parked with a snow mountain view behind them. VQA What numbers are displayed on the front of the bus on the right? Short answer: Assistant: Response 6044 Multimodal Dialogue What might be the purpose of the buses in this location? Assistant: Response It is plausible that they are in this location for multiple reasons. Some possible reasons might be: 1)…2)… Table 11: Examples of different instructions with different output formats. We use bold fonts to indicate different instructions.
Appendix G Training Efficiency
Table 12 summarizes the key factors in training EasyGen. The training process of EasyGen is computationally efficient, especially with the utilization of the parameter-efficient fine-tuning method LoRa (Hu et al., 2021). To enable multimodal response generation, we further train the aligned EasyGen. This process entails fine-tuning the LLM (FlanT5XL) on the PhotoChat dataset for 2 epochs, which typically requires approximately 4 A100 (80G) GPU hours.
Appendix H Impact of Alignment Manners
In Table 13, we investigate the impact of different alignment manners on EasyGen. After removing the ITDM loss, the performance is slightly weaker than the original model. It is evident that the MSE Loss can help to align the semantic spaces of the two models. Furthermore, the performance of the model will drop significantly after removing the cross-entropy loss, suggesting that constraints via the language model play a key role.
Model Trainable Param. Training Images Training Cost
Pre-training
 BiDiffuser 952M 169K 120 (A100 80GB) GPU hours
Alignment
 Projection Layers + [Uncaptioned image] T5XL 4M 163K 20 (RTX3090 24GB) GPU hours
 Projection Layers + [Uncaptioned image] T5XL 3B 173K 20 (A100 80GB) GPU hours
 Projection Layers + [Uncaptioned image] Vicuna 7B 7B 173K 72 (A100 80GB) GPU hours
 Projection Layers + [Uncaptioned image] Vicuna 7B(LoRa) 610M 173K 20 (A100 80GB) GPU hours
Table 12: EasyGen’s trainable parameters, training data size, and training cost during alignment process.
Model NoCaps (val) COCO (Karpathy) OK-VQA GQA
CIDEr SPICE SPICE BLEU@4 CIDEr Accuracy Accuracy
EasyGen Mid-Align FlanT5XL 121.2 15.5 25.1 43.5 145.7 31.5 22.6
+ w/o ITDM 118.6 15.3 24.8 42.2 141.5 - -
+ w/o ITG 93.2 12.9 23.0 35.1 127.6 - -
Table 13: Ablation studies on the instruction-tuning process and loss functions.
Figure 7: Examples of the generated responses on PhotoChat dataset. The text highlighted in red indicates the objects present in the image. The turns prefixed with A/B denote the given context.
Appendix I More Qualitative Results
We present several instances on PhotoChat dataset in Figure 7 and the image-captioning task in Figure 8. In Figure 9, 11, 10, we compare EasyGen with state-of-the-art multimodal language models. The responses of MiniGPT-4, LLaVA, mPLUG-owl and InstructBlip are obtained from their official demos. Morever, in Figure 12, 13, we show EasyGen’s ability to accept multimodal inputs and generate multimodal responses.
Figure 8: Examples of image captioning results by EasyGen.
Figure 9: In this case study, for the first question, EasyGen can give an accurate answer including the background information of the image. With the image generation ability of BiDiffuser, EasyGen can generate visual responses.
Figure 10: In this case study, for the first question, EasyGen can give an accurate answer, but the responses of the other two models are a bit biased. For the second question, EasyGen and LLaVA both give reasonable advice.
Figure 11: From this example, we can find that the response from EasyGen is more comprehensive and coherent. This shows EasyGen can give reasonable suggestions based on the given image.
Figure 12: Example of multimodal response generation.
Figure 13: Example of multimodal response generation.